is 3x-4=12 an linear epuatoin
Answer by stanbon(75887)

is 3x-4=12 an linear epuatoin
----------------------
No because it simplifies to x = 16/3 which is not a function.
-------------
x = 16/3 cannot be written in the form ax+by=c where
b is not zero.
---------------------
Cheers,
Stan H.